OpenStudy (anonymous):

The population of the United States is approximately 3x10^6 people. The population of Germany is approximately 8x10^7. Which statement about the populations of these countries is true ?

imqwerty (imqwerty):

what are the statements?

OpenStudy (astrophysics):

I'm going to guess you have to make a ratio

imqwerty (imqwerty):

my guess- which one is more or less

OpenStudy (anonymous):

A) The population of Germany is almost 3 times the population of the U.S B) The population of Germany is almost 4 times the population of the U.S C) The population of the U.S is almost 3 times the population of of Germany D) The population of the U.S is almost 4 times the population of Germany

imqwerty (imqwerty):

aww

OpenStudy (astrophysics):

lolol

imqwerty (imqwerty):

check your question that must be population of germany= \(8 \times 10^{\color{red}6}\)

OpenStudy (astrophysics):

Yeah 10^7 is way too much

OpenStudy (astrophysics):

for the question

imqwerty (imqwerty):

lol yeah

OpenStudy (anonymous):

No it's 8x10^7

OpenStudy (astrophysics):

3*10^7?

OpenStudy (astrophysics):

I think it's 3*10^7 in any case you can set a ratio up haha

imqwerty (imqwerty):

check your options

OpenStudy (astrophysics):

\[\frac{ 8 \times 10 ^7 }{3 \times 10^7 } = k\] where k will give you how much more people Germany will have than the US

OpenStudy (astrophysics):

Either they're both to the power of 6 or 7 because otherwise your options are wrong

OpenStudy (anonymous):

Yeah, I guess my options are wrong

imqwerty (imqwerty):

\(\Huge\ddot\smile\)

OpenStudy (astrophysics):

Haha and yeah, the links qwerty gave are the real populations

OpenStudy (anonymous):

yeah lol

OpenStudy (astrophysics):

Oh you know what...

OpenStudy (astrophysics):

They have the populations backwards, I think they meant \[3 \times 10^7\] for US and \[8 \times 10^ 6\] for Germany, that's pretty close to real populations and then if you do \[\frac{ 3 \times 10^7 }{ 8 \times 10^6 } = k \] you will get one of your options

imqwerty (imqwerty):

yeah this seems correct cuz germany's population has always been less than US's

OpenStudy (anonymous):

I got 0.375

OpenStudy (astrophysics):

Try again

OpenStudy (anonymous):

2.6 ?

OpenStudy (astrophysics):

Why I did it for you, plug in those numbers

imqwerty (imqwerty):

\(\huge \frac{ 3 \times 10^7 }{ 8 \times 10^6 } = k\)

OpenStudy (anonymous):

okay so I did \[3\div8 \] and I got 0.375

OpenStudy (astrophysics):

Put the powers..

OpenStudy (anonymous):

I totally forgot my bad

OpenStudy (anonymous):

3.7

OpenStudy (astrophysics):

Yes, 3.75 approximately 4

OpenStudy (anonymous):

oh yeah

OpenStudy (astrophysics):

What does that imply?

OpenStudy (anonymous):

so the answer is D

OpenStudy (astrophysics):

Yes

imqwerty (imqwerty):

:)

OpenStudy (anonymous):

THANK YOU SO MUCH !!!!!!!!!!!!11

imqwerty (imqwerty):

\(\huge\cal\color{blue}{ヾ(⌐■‿■)ノ♪}\color{pink}{*:・゚}\color{gold}{✧*:}\color{red}{・゚✧*}\color{green}{:・゚:・゚}\color{violet}{✧*:・゚✧*:・゚✧*:・゚✧*:・゚✧*:・゚✧}\)

OpenStudy (anonymous):

i'm here! XD